OpenStudy (anonymous):

calculate the average speed for a round trip with a going average speed of 48 and a returning average speed of 36.9 miles per hour

OpenStudy (phi):

the average speed is the total distance divided by the total time use speed * time = distance or, for our case time = distance/speed Going, it took time T1 (which we don't know). We went a distance D (we also don't know) But we do know the speed was 48 T1 = D/48 Returning, using the same formula, but a different speed 36.9, we find the time to return which we will call T2: T2 = D/36.9

OpenStudy (phi):

The total distance there and back is D + D or 2D total time is T1+ T2 the average speed is \[ Speed_{avg}= \frac{2D}{T_1+T_2} \] can you finish?

OpenStudy (anonymous):

distance for both is 240 miles time for the 48 mph is 5 hours and the time for 36.9 mph is 6.5 or 6 hours and 30 minutes. For some reason I really dont understand this problem every other one had been easy

OpenStudy (phi):

Did they give you the distance ?

OpenStudy (anonymous):

240 miles going and coming back

OpenStudy (phi):

Forget about the speed going or coming back. That does not matter what matters is total distance divided by total time

OpenStudy (anonymous):

probably not this but how you said it like this? 240+240 = 480 5+6.5= 11.5 and 480 divided by 11.5?

OpenStudy (phi):

exactly

OpenStudy (anonymous):

Ok, yah thank you

OpenStudy (phi):

these problems are confusing because you probably want to average the speed going and returning, and that is not right. And it's hard to see it's not right. But you can remember this: speed is *always* distance divided by time so if you can figure out the distance and the time it took, you can find the speed.

OpenStudy (anonymous):

yah ok thanks